OpenStudy (anonymous):

what is the change in enthalpy for the following reaction? 2H2(g) + 2C(s) + O2(g) d C2H4OH(l)? C2H5OH(l) + 2O2(g) -> 2CO2(g) + 2H2O(l),   DH = -875 kJ C(s) + O2(g) -> CO2(g),  DH = -394.51 kJ H2(g) + ½O2(g) -> H2O(l),  DH = -285.8 kJ is it -194.7 kJ...?

OpenStudy (chmvijay):

nooo it will be -477 KJ

OpenStudy (anonymous):

could you explain how?

OpenStudy (chmvijay):

reverse the equation 2CO2(g) + 2H2O(l)===> C2H5OH(l) + 2O2(g) DH = +875 kJ -1 multiply next equations by 2 2C(s) + 2O2(g) -> 2CO2(g), DH = -394.51*2 kJ -2 2H2(g) + O2(g) -> 2H2O(l), DH = -285.8 8*2 kJ -3 add all the three you get 2H2(g) + 2C(s) + O2(g) d C2H4OH(l) =-789-571+875 = -477 KJ hope u got it

OpenStudy (anonymous):

but wait, would that be -486 kJ instead?

OpenStudy (anonymous):

wouldn't*

OpenStudy (chmvijay):

yaa ur right its 485.8 something :)

OpenStudy (anonymous):

okay!(: Thank you. I truly appreciate your help.

OpenStudy (chmvijay):

you are welcome :)

OpenStudy (anonymous):

could you help me with another one? If so, do I need to post in a different thingy or can I just ask in here?

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

you just post in another :)

OpenStudy (anonymous):

do I have to 'close' this one?